A 8.00L tank at −2.4°C is filled with 14.4g of sulfur hexafluoride gas and 20.0g of...

A

8.00L

tank at

−2.4°C

is filled with

14.4g

of sulfur hexafluoride gas and

20.0g

of sulfur tetrafluoride gas. You can assume both gases behave as ideal gases under these conditions.

Calculate the mole fraction of each gas. Be sure each of your answer entries has the correct number of significant digits.
Answer #1

Molar mass of SF6,

MM = 1*MM(S) + 6*MM(F)

= 1*32.07 + 6*19.0

= 146.07 g/mol

Molar mass of SF4,

MM = 1*MM(S) + 4*MM(F)

= 1*32.07 + 4*19.0

= 108.07 g/mol

n(SF6) = mass/molar mass

= 14.4/146.07

= 0.0986

n(SF4) = mass/molar mass

= 20.0/108.07

= 0.1851

n(SF6),n1 = 0.0986 mol

n(SF4),n2 = 0.1851 mol

Total number of mol = n1+n2

= 0.0986 + 0.1851

= 0.2836 mol

Mole fraction of each components are

X(SF6) = n1/total mol

= 0.0986/0.2836

= 0.3476

X(SF4) = n2/total mol

= 0.1851/0.2836

= 0.6524

Answer:

X(SF6) = 0.348

X(SF4) = 0.652